Two angles are supplementary. One is 15 degrees more than four times the other. Find the measures of the angles.
What is the smaller angle?
What is the other angle?

Answer by nerdybill(7384)

Two angles are supplementary. One is 15 degrees more than four times the other. Find the measures of the angles.
What is the smaller angle?
What is the other angle?
.
Let x = smaller angle
then
4x+15 = other angle
.
x + 4x+15 = 180 (definition of supplementary)
5x+15 = 180
5x = 165
x = 33 degrees (smaller angle)
.
Other angle:
4x+15 = 4(33)+15 = 132+15 = 147 degrees